# font=sf,footnotesize UN, Nantes]Nantes Université, CNRS, CEISAM UMR 6230, F-44000 Nantes, France

### v1

@@ 3.3.5 BSE/ vs TBE
Let us now discuss BSE/. The plots displayed in Figure 8 and the data of Table S29 indicate that HF starting quantities are not recommended: they lead to strong overestimations of the VTEs with MSEs larger than 0.3 eV irrespective of the applications of the TDA and/or the selection of ev or . Large errors are obtained for all subsets of ESs. Although they lead to negative MSEs, KS orbitals are clearly more adequate for the present systems. Globally, the more rudimental scheme is detrimental to the final accuracy, especially with PBE0 since the VTEs are largely underestimated, consistent with earlier benchmarks performed on Thiel’s set. 194, 190, 28 requires XCFs with a large share of EXX as, for example, CAM-B3LYP. 190 The smallest MAE and SDE, both of 0.21 eV, are reached with TDA-BSE/ev@CAM-B3LYP. Such deviations can likely be viewed as rather large since several XCFs deliver more accurate results when used in TD-DFT calculations.
@@ 3.4 Further discussion
Another aspect that we already briefly evoked in Section 3.3.4 is the importance of accurate references for benchmarking purposes. This central point is illustrated in Figure 9 where we provide the MAEs obtained at the TDA-DFT level considering the entire set of XCFs using the TBE, CCSD, or ADC(2) values as references. We underline that all ESs are present for all methods (both reference and tested approaches), allowing well-grounded comparisons. Obviously, benchmarking TDA-DFT with CCSD references is not a fantastic idea: the MAE of the GHs with a low EXX percentage significantly increases (e.g., B3LYP’s MAE goes from 0.31 to 0.45 eV), whereas the MAE of RSHs with a large attenuation parameter is strongly underestimated (e.g., B97X’s MAE goes from 0.32 to 0.17 eV). This is the logical consequence of the tendency of CCSD to overestimate the VTEs, as illustrated in Figure 2. In other words, CCSD favors XCFs with a large amount of EXX, since they exhibit the same overshooting predisposition.
From the wave function results of Table 2, one could argue that opting for ADC(2) as a reference method is a better choice since it presents a more balanced error pattern with a near-zero MSE of eV (Table 2). ADC(2) is actually the reference method of the very recent DELFI “oracle” whose aim is to predict the most suitable XCF for a given compound through a machine-learning-based approach. 206 From Figure 9, one indeed notices quite similar MAE patterns for most XCFs when using the TBE or ADC(2) values as references. However, by taking a closer look, one sees that the MAEs predicted for the most accurate XCFs are significantly overshot when one relies on ADC(2). For example, SOS-B88PP86’s and SOS-PBEPP86’s MAEs are shifted from 0.12 to 0.19 eV and from 0.11 to 0.18 eV, respectively when replacing the TBEs by the ADC(2) data. Actually, ADC(2) would also incorrectly foresee B2PLYP as the best performing DH with a MAE of 0.13 eV. In other words, ADC(2) can probably be employed to pinpoint poor functionals, but appears not accurate enough to highlight the best ones.
@@ 4. Conclusions
With these reference values at hand, we benchmarked more computationally affordable methods. We first considered 14 second- and third-order wave function methods. A key finding is that the predictions of CCSD are poorer than those of CC2 and ADC(2) in terms of absolute errors (which was expected) and standard deviations (which differs from results obtained for smaller compounds). The conclusion that CC2 is less consistent than CCSD does not hold for larger, real-life molecules. While some might view this finding as good news for practical applications, as CC2 is a relatively cheap method, our data suggest that significant improvements might require much more expensive wave function approaches like CC3 or CCSDT. Correcting CCSD VTEs with perturbative estimates of the triples still does not outperform CC2 in terms of absolute deviation. Consistent with the results obtained in the original QUEST, we found that the ADC(2.5) protocol, which averages ADC(2) and ADC(3) VTEs, is the best performing method among all and wave function methods we tested here, although it suffers from a small, nearly systematic underestimation trend.
We have also considered a large set of density-based models, including many global, range-separated, local, and double hybrid XCF, applied with or without the TDA, and performed calculations in both the TD-DFT and BSE/ contexts. These methods are significantly less computationally demanding than their wave function counterparts and are typically used when modeling organic dyes. Gratifyingly, our conclusions, both qualitatively and quantitatively, align well with the most recent benchmarks.

### v2

@@ 3.3.5 BSE/ vs TBE
Let us now discuss BSE/. The plots displayed in Figure 8 and the data of Table S29 indicate that HF starting quantities are not recommended: they lead to strong overestimations of the VTEs with MSEs larger than 0.3 eV irrespective of the applications of the TDA and/or the selection of ev or . Large errors are obtained for all subsets of ESs. Although they lead to negative MSEs, KS orbitals are clearly more adequate for the present systems. Globally, the more rudimental scheme is detrimental to the final accuracy, especially with PBE0 since the VTEs are largely underestimated, consistent with earlier benchmarks performed on Thiel’s set. 194, 190, 28 requires XCFs with a large share of EXX as, for example, CAM-B3LYP. 190 The smallest MAE and SDE, both of 0.20 eV, are reached with TDA-BSE/ev@CAM-B3LYP. Such deviations can likely be viewed as rather large since several XCFs deliver more accurate results when used in TD-DFT calculations.
@@ 3.4 Further discussion
Another aspect that we already briefly evoked in Section 3.3.4 is the importance of accurate references for benchmarking purposes. This central point is illustrated in Figure 9 where we provide the MAEs obtained at the TDA-DFT level considering the entire set of XCFs using the TBE, CCSD, or ADC(2) values as references. We underline that all ESs are present for all methods (both reference and tested approaches), allowing well-grounded comparisons. Obviously, benchmarking TDA-DFT with CCSD references is not a fantastic idea: the MAE of the GHs with a low EXX percentage significantly increases (e.g., B3LYP’s MAE goes from 0.31 to 0.45 eV), whereas the MAE of RSHs with a large attenuation parameter is strongly underestimated (e.g., B97X’s MAE goes from 0.30 to 0.17 eV). This is the logical consequence of the tendency of CCSD to overestimate the VTEs, as illustrated in Figure 2. In other words, CCSD favors XCFs with a large amount of EXX, since they exhibit the same overshooting predisposition.
From the wave function results of Table 2, one could argue that opting for ADC(2) as a reference method is a better choice since it presents a more balanced error pattern with a near-zero MSE of eV (Table 2). ADC(2) is actually the reference method of the very recent DELFI “oracle” whose aim is to predict the most suitable XCF for a given compound through a machine-learning-based approach. 206 From Figure 9, one indeed notices quite similar MAE patterns for most XCFs when using the TBE or ADC(2) values as references. However, by taking a closer look, one sees that the MAEs predicted for the most accurate XCFs are significantly overshot when one relies on ADC(2). For example, SOS-B88PP86’s and SOS-PBEPP86’s MAEs are shifted from 0.12 to 0.19 eV and from 0.10 to 0.18 eV, respectively when replacing the TBEs by the ADC(2) data. Actually, ADC(2) would also incorrectly foresee B2PLYP as the best performing DH with a MAE of 0.13 eV. In other words, ADC(2) can probably be employed to pinpoint poor functionals, but appears not accurate enough to highlight the best ones.
@@ 4. Conclusions
With these reference values at hand, we benchmarked more computationally affordable methods. Indeed, CC3 is generally too expensive to be used in practice. To give an order of magnitude, computing the four CC3/AVTZ VTEs of aza-BODIPY took us 29 days of (human) time on a 2Tb/48 cores node, though such timing of course depends on the details of both the implementation and architecture. In our benchmark, we first considered 14 second- and third-order wave function methods. A key finding is that the predictions of CCSD are poorer than those of CC2 and ADC(2) in terms of absolute errors (which was expected) and standard deviations (which differs from results obtained for smaller compounds). The conclusion that CC2 is less consistent than CCSD does not hold for larger, real-life molecules. While some might view this finding as good news for practical applications, as CC2 is a relatively cheap method (e.g., the same four ESs of aza-BODIPY could be computed in less than 1h on a standard 128Gb node), our data suggest that significant improvements might require much more expensive wave function approaches like CC3 or CCSDT. Correcting CCSD VTEs with perturbative estimates of the triples still does not outperform CC2 in terms of absolute deviation. Consistent with the results obtained in the original QUEST, we found that the ADC(2.5) protocol, which averages ADC(2) and ADC(3) VTEs, is the best performing method among all and wave function methods we tested here, although it suffers from a small, nearly systematic underestimation trend. In practice however, the ADC(3) calculations will be the bottleneck in the ADC(2.5) protocol.
We have also considered a large set of computationally more effective models, including TD-DFT with many global, range-separated, local, and double hybrid XCF, applied with or without the TDA, as well as various BSE/ models. These methods typically require a fraction of the computational effort of their wave function counterparts and are typically used when modeling organic dyes. Gratifyingly, our conclusions for both TD-DFT and BSE/, both qualitatively and quantitatively, align well with the most recent benchmarks.